# Register Any Point: Scaling 3D Point Cloud Registration by Flow Matching

### v1

@@ Appendix B Details on the Selected Baselines
PARENet [yao2024iccv-parenet] Deep learning ECCV’24 Position-Aware Rotation-Equivariant Network designed for robust registration in low-overlap scenarios by learning rotation-equivariant features and position-aware representations that handle partial point clouds effectively
BUFFER-X [seo2025iccv] Deep learning ICCV’25 Extended BUFFER method with improved zero-shot generalization capabilities, adaptively determining voxel sizes and search radii, bypassing learned keypoint detectors using farthest point sampling, and employing patch-wise scale normalization for consistent performance across diverse scenes
Multi-view Registration Baselines (all follow the pairwise registration and pose graph optimization two-stage pipeline)
FGR [zhou2016eccv] + PGO Conventional ECCV’16 Fast Global Registration performs pairwise alignment between all point cloud pairs. Then we apply grow-based algorithm to get the initial guess of the scan poses, followed by pose graph optimization (PGO) implemented by Open3D to enforce global consistency across all views
BUFFER-X [seo2025iccv] + PGO Deep learning ICCV’25 We use BUFFER-X for pairwise registration and apply grow-based algorithm to get the initial guess of the scan poses. We then apply PGO implemented by Open3D to achieve globally consistent multi-view alignment
LMVR [gojcic2020cvpr] Deep learning CVPR’20 Learned Multi-View Registration end-to-end network that jointly estimates transformations for multiple point clouds by learning to predict the exhaustive pairwise transformations and optimizing them simultaneously with a transformation synchronization step in a unified framework
SGHR [wang2023cvpr-sghr] Deep learning CVPR’23 Uses NetVLAD-based global features to estimate overlaps and construct a sparse pose graph, reducing the number of pairwise registrations. Applies history reweighting in IRLS that considers previous iterations’ poses to prevent outlier domination and ensure stable convergence
Appendix C Additional Experimental Results
Comprehensive results on all validation and testing datasets
To provide a holistic view of our model’s performance across all validation datasets used during training, we summarize the multi-view registration metrics—registration success rate (SR), Chamfer Distance (CD), mean rotation error (RE), and mean translation error (TE)—for both average ($S=1$) and rigidity-selected ($S=3$) inference settings on both scan- and submap-level data in Tab. 8. As mentioned in the main paper, $S$ is the number of random generations for inference. We show both the results for validation samples in both the seen sequences and the unseen sequences. Overall, our model has better performance on the scan-level samples than the submap-level samples, and better performance on the seen sequences than on the unseen sequences. Besides, the rigidity-selected inference setting generally outperforms the average inference setting, demonstrating the effectiveness of our rigidity-based selection strategy.
We also provide the comprehensive experimental results on the testing datasets (including those TLS datasets) in Tab. 9.
Table 8: Comprehensive multi-view registration performance on the validation sets of all datasets. We report registration success rate (SR, in %), Chamfer Distance (CD, in meters), mean rotation error (RE, in degrees), and mean translation error (TE, in meters) for both average ($S=1$) and rigidity-selected ($S=3$) inference settings on both scan- and submap-level data where applicable. We show the results of validation samples on both the seen and unseen sequences.
Average ($S=1$) Rigidity Selected ($S=3$)
Dataset # Views Unseen Sequence SR [%] $\uparrow$ CD [m] $\downarrow$ RE [] $\downarrow$ TE [m] $\downarrow$ SR [%] $\uparrow$ CD [m] $\downarrow$ RE [] $\downarrow$ TE [m] $\downarrow$
Outdoor LiDAR
KITTI Scan [2,8] Yes 82.00 1.46 15.84 8.65 83.11 1.38 14.41 8.09
No 91.60 0.73 3.72 3.22 91.87 0.82 4.29 3.81
KITTI Submap [2,10] Yes 32.89 12.77 35.18 74.45 33.85 12.67 34.69 73.93
No 51.56 3.22 5.22 11.59 54.52 2.89 4.66 9.81
KITTI360 Scan [2,8] Yes 85.04 0.76 4.57 7.17 86.28 0.72 4.28 7.05
No 96.25 0.33 0.75 0.81 97.35 0.31 0.69 0.75
KITTI360 Submap [2,10] Yes 39.56 10.31 34.68 57.36 41.53 10.19 34.03 56.77
No 72.90 2.27 5.15 10.64 74.27 2.21 5.16 10.43
Apollo Scan [2,8] Yes 73.23 1.13 8.78 14.00 73.36 1.13 8.73 13.96
No 90.26 0.64 2.50 3.42 91.65 0.52 2.12 2.84
Apollo Submap [2,10] Yes 91.88 1.57 3.45 9.84 92.15 1.65 3.51 10.19
No 94.12 0.66 1.42 2.30 95.13 0.61 1.42 1.97
MulRAN Scan [2,5] No 86.30 0.97 6.41 4.74 88.89 0.86 5.78 3.67
MulRAN Submap [2,10] Yes 69.20 7.08 12.71 38.36 69.93 6.71 12.74 37.61
No 66.43 2.45 5.30 10.85 70.51 2.19 5.12 10.25
Oxford Spires Scan [2,10] Yes 27.80 2.49 26.18 10.47 29.18 2.51 25.28 10.36
No 73.77 0.82 4.07 1.59 75.00 0.77 3.63 1.41
Oxford Spires Submap [2,10] Yes 85.09 0.74 3.80 1.39 89.47 0.66 2.86 1.20
No 82.42 1.27 4.63 2.37 81.82 1.23 4.94 2.31
VBR Scan [2,8] Yes 40.28 2.53 20.43 16.88 40.74 2.56 20.94 16.93
No 81.76 0.67 2.47 2.25 83.43 0.64 2.27 2.09
VBR Submap [2,10] Yes 15.91 10.95 39.43 48.23 16.82 10.86 38.81 47.77
No 51.83 2.83 7.27 10.28 53.92 2.65 6.78 9.75
UrbanNav Scan [2,8] Yes 60.00 3.39 27.32 14.42 59.43 3.36 27.88 15.04
No 80.91 1.10 5.17 6.29 81.26 1.01 5.51 6.02
UrbanNav Submap [2,10] Yes 32.84 6.42 31.12 33.49 32.11 6.37 32.05 35.10
No 80.28 1.58 4.91 5.74 77.56 1.60 4.63 5.85
HeLiPR Scan [2,8] Yes 88.89 0.55 2.38 4.93 88.62 0.54 2.49 4.78
HeLiPR Submap [2,10] No 92.31 0.86 1.85 2.18 92.31 0.94 1.80 2.44
Boreas Scan [2,5] No 97.68 0.32 2.55 2.05 97.39 0.25 2.40 1.79
Boreas Submap [2,8] No 96.04 0.81 1.59 4.94 96.52 0.88 1.50 4.37
WildPlace Submap [2,5] No 86.64 0.60 3.87 4.41 86.35 0.62 4.36 4.83
NuScenes Pairwise Scan [2,2] No 97.62 0.28 1.47 0.79 97.86 0.27 1.42 0.73
KITTI-Carla Submap [2,10] No 56.49 2.69 5.20 9.13 59.08 2.56 4.66 9.81
Indoor Depth Camera
3DMatch Scan [2,2] Yes 76.17 0.09 16.34 0.46 76.57 0.09 16.05 0.45
No 88.33 0.05 6.02 0.16 88.66 0.05 5.53 0.16
3DMatch Submap [2,16] Yes 30.61 0.34 38.12 1.46 31.64 0.34 38.96 1.46
No 42.36 0.27 25.13 1.15 43.25 0.27 24.61 1.15
ScanNet Submap [2,16] Yes 66.22 0.13 10.90 0.35 71.10 0.12 9.59 0.33
No 80.27 0.09 11.00 0.31 81.87 0.08 10.53 0.29
ScanNet++ Scan [2,12] Yes 74.53 0.10 14.28 0.34 74.83 0.10 14.35 0.33
No 83.18 0.08 5.11 0.23 84.02 0.07 5.30 0.23
NSS Pairwise Scan [2,2] No 96.21 0.03 2.78 0.12 96.35 0.02 2.61 0.11
NSS Multi-view Scan [2,20] Yes 53.98 0.30 19.33 1.41 54.14 0.30 19.24 1.41
Object-centric
ModelNet-40 [2,2] Yes 97.18 0.006 4.81 0.016 97.33 0.006 4.65 0.015
Table 9: Comprehensive multi-view registration performance on the testing datasets. We report registration success rate (SR, in %), Chamfer Distance (CD, in meters), mean rotation error (RE, in degrees), and mean translation error (TE, in meters) for both average ($S=1$) and rigidity-selected ($S=5$) inference settings on both scan- and submap-level data where applicable.
Average ($S=1$) Rigidity Selected ($S=5$)
Dataset # Views Type Zero-shot SR [%] $\uparrow$ CD [m] $\downarrow$ RE [] $\downarrow$ TE [m] $\downarrow$ SR [%] $\uparrow$ CD [m] $\downarrow$ RE [] $\downarrow$ TE [m] $\downarrow$
Pairwise Registration
3DMatch [zeng2017cvpr] [2,2] Scan No 95.27 0.02 3.76 0.08 95.64 0.02 3.71 0.07
3DLoMatch [huang2021cvpr-predator] [2,2] Scan No 77.43 0.08 13.46 0.36 78.22 0.08 13.22 0.34
NSS [sun2025jprs] [2,2] Scan No 96.39 0.03 2.83 0.15 96.18 0.03 2.78 0.15
TIERS [sier2023rs-tiers] [2,2] Scan Yes 94.75 0.13 3.87 0.37 95.09 0.12 3.53 0.34
KITTI [geiger2012cvpr] [2,2] Scan No 100.00 0.14 0.20 0.09 100.00 0.13 0.20 0.09
Waymo [sun2020cvpr] [2,2] Scan Yes 98.51 0.22 2.69 0.69 99.47 0.18 2.33 0.45
KAIST [jung2024ijrr] [2,2] Scan Yes 99.05 0.19 1.45 0.44 99.15 0.18 1.48 0.46
MIT [tian2023iros-kimeramultidata] [2,2] Scan Yes 98.96 0.28 1.98 0.43 99.13 0.24 1.93 0.26
ETH [pomerleau2012ijrr] [2,2] Scan Yes 98.30 0.07 2.07 0.15 98.04 0.06 1.97 0.13
NCD [ramezani2020iros] [2,2] Scan Yes 99.34 0.32 2.14 0.96 99.34 0.24 1.65 0.55
Multi-view Registration
3DMatch [zeng2017cvpr] [3,12] Scan No 75.29 0.11 7.46 0.23 76.50 0.10 7.30 0.22
ScanNet [dai2017cvpr] [3,12] Scan No 63.83 0.13 16.54 0.42 64.04 0.12 13.66 0.35
Waymo [sun2020cvpr] [3,12] Scan Yes 79.17 0.51 1.38 2.12 87.50 0.23 1.10 1.56
KAIST [jung2024ijrr] [3,8] Submap Yes 88.12 3.06 5.37 9.14 87.40 2.97 5.52 9.36
WHU-TLS [dong2020jprs-whutls] [5,10] Scan Yes 34.34 6.34 24.55 46.50 34.31 6.08 23.05 44.95
ETH-TLS [theiler2015jprs] [5,8] Scan Yes 60.00 0.86 6.34 3.48 75.00 0.71 4.37 2.39
Additional ablation studies
For additional ablation studies, we provide detailed results in Tab. 10, as an extension of the ablation studies in Table 4 from the main paper.
It is shown that with more generation steps, the performance of our model generally improves. However, the improvement becomes marginal beyond 20 steps. Besides, using only indoor training data [K] also degrades the performance, even on the indoor datasets. We additionally provide the results for the best-of-3, best-of-5, and best-of-10 generation in Tab. 10 to show the upper bound performance of our model at different stochastic budget.
Table 10: Extended ablation studies. We report the registration success rate (%) on datasets for both pairwise and multi-view registration tasks. Best results are shown in bold, second best are underscored.
Setting Pairwise Multi-view
3DMatch 3DLoMatch ScanNet KAIST
[A] Ours (S = 5) 95.64 78.22 64.04 87.40
Inference
[B] w/o rigidity forcing 94.65 74.63 59.76 83.69
[C] w/o rigidity-based selection 95.27 77.43 63.83 88.12
[D] w/ 1-step generation 90.35 61.05 48.33 52.95
[E] w/ 5-step generation 95.52 76.83 58.91 87.91
[F] w/ 20-step generation 96.07 79.30 66.61 91.04
[G] w/ 50-step generation 95.88 79.45 65.37 91.46
Model
[H] w/o local feature extraction 83.28 50.05 21.82 47.20
[I] w/ $L=6$ transformer blocks 90.90 57.59 45.71 73.77
[J] w/ $L=8$ transformer blocks 93.24 71.63 56.24 85.14
[K] w/ only indoor training data 92.19 70.02 57.40 12.25
Stochastic
[L] Best-of-3 generation 95.45 79.24 65.96 91.12
[M] Best-of-5 generation 95.88 81.09 69.81 92.59
[N] Best-of-10 generation 96.31 81.91 72.38 93.34
Appendix D Additional Qualitative Results
We provide additional qualitative results of our model on the pairwise and multi-view point cloud registration tasks on various datasets in this section, as shown in Fig. 7, Fig. 8, Fig. 9, and Fig. 10.
We also provide some failure cases of our model in Fig. 11. It is shown that our model sometimes fails to correctly register the point clouds in the cases of (i) very low overlap ratio or even no overlap due to the thickness of walls (NSS, ScanNet++); (ii) complex scenes unseen during training (TLS tunnel, TLS forest). From the examples shown for NSS, it is shown that our model can still guess the reasonable global layout of the scene for the ambiguous failed cases.
Figure 7: Qualitative results of our model on the pairwise registration benchmarks with large scan interval (Waymo, KITTI, KAIST, NCD, MIT, TIERS) or low overlap ratio (NSS, 3DLoMatch).

### v2

@@ Appendix C Details on the Selected Baselines
PARENet [yao2024iccv-parenet] Deep learning ECCV’24 Rotation-equivariant, position-aware network for robust registration in low-overlap scenarios. Trained on 3DMatch or KITTI.
BUFFER-X [seo2025iccv] Deep learning ICCV’25 Extends BUFFER with improved zero-shot generalization via adaptive voxel sizing, farthest-point sampling, and patch-wise scale normalization. Originally trained on 3DMatch or KITTI; we additionally retrain a variant, denoted BUFFER-X, on our training data.
Multi-view Registration Baselines
Pairwise registration + PGO [choi2015cvpr] - CVPR’15 All-pair pairwise registration followed by a grow-based initialization and pose graph optimization, implemented by Open3D [zhou2018arxiv].
SGHR [wang2023cvpr-sghr] Deep learning CVPR’23 Constructs a sparse pose graph using NetVLAD-based overlap scores and conducts pairwise registration with YOHO [wang2022mm-yoho] on the edges from the sparse pose graph. Then it applies history-reweighted iterative reweighted least squares (IRLS) for stable, outlier-robust convergence. Trained on 3DMatch.
RPF [sun2025neurips] Deep learning NeurIPS’25 A flow-matching framework in Euclidean space that formulates multi-part shape assembly as a conditional generative problem by learning a continuous point-wise velocity field that transports points to their assembled state. Trained on ModelNet and several object-centric shape assembly datasets.
Appendix D Additional Experimental Results
D.1 Additional results on the pairwise registration with low overlap
To further demonstrate our model’s robustness to low overlap between point clouds, especially in the outdoor LiDAR scenarios, we follow EYOC [liu2024cvpr-eyoc] to curate testing data on both the KITTI [geiger2012cvpr] and Waymo [sun2020cvpr] with increasing spatial distance from 10 m to 50 m between point clouds (thus decreasing overlap). Note that KITTI at 10 m is the setting used in the standard pairwise registration benchmark. As shown in Fig.˜6, our model shows superior performance over state-of-the-art methods such as BUFFER-X [seo2025iccv], Predator [huang2021cvpr-predator], FCGF [choy2019iccv], and EYOC [liu2024cvpr-eyoc] with increasing scan distance.
D.2 Additional results on indoor multi-view registration benchmark
For multi-view registration evaluation, we also follow prior work [gojcic2020cvpr, wang2023cvpr-sghr] and use the common multi-view registration benchmarks on 3DMatch [zeng2017cvpr] and ScanNet [dai2017cvpr]. They are both indoor depth camera datasets. We evaluate on sparse-view subsets (with $3\leq N\leq 12$) of 3DMatch and ScanNet and report the translation error (TE), rotation error (RE), and Chamfer distance (CD). Comparisons with the baseline methods are shown in Tab.˜7, demonstrating superior performance. The sparse-view setting makes the pairwise registration more challenging, thus most of the two-stage multi-view registration baselines fail to achieve good performance.
Table 7: Quantitative comparison of multi-view point cloud registration on 3DMatch and ScanNet dataset under the sparse view setting (with view count $3\leq N\leq 12$). We report the mean rotation error (RE) in degrees, mean translation error (TE) in meters, and Chamfer distance (CD) in meters. Best results are shown in bold.
3DMatch ScanNet
Method RE $\downarrow$ TE $\downarrow$ CD $\downarrow$ RE $\downarrow$ TE $\downarrow$ CD $\downarrow$
FGR [zhou2016eccv] + PGO 52.81 0.71 0.49 68.80 1.43 0.76
BUFFER-X [seo2025iccv] + PGO 48.16 0.74 0.48 47.63 1.31 0.52
LMVR [gojcic2020cvpr] 15.46 0.44 0.16 46.10 0.87 0.50
SGHR [wang2023cvpr-sghr] 50.28 0.78 0.53 23.59 0.64 0.34
RAP (Ours) 7.27 0.23 0.11 13.85 0.34 0.12
D.3 Additional results on cross-domain multi-view registration benchmark
We show the comparison results on the cross-domain multi-view registration benchmark in Tab.˜8, where we use the strict registration success rate (with a threshold of 0.5% for the normalized translation error and 3 for the rotation error) instead of the standard threshold (2.5% for the normalized translation error and 15 for the rotation error) used in the main paper.
Table 8: Comparison of the zero-shot testing performance for multi-view registration on the cross-domain multi-view registration benchmark. We report the strict registration success rate (%) calculated for edges and graphs (with a threshold of 0.5% for the normalized translation error and 3 for the rotation error) in different scenarios as well as the average runtime. The best result is in bold, and the second best is underscored.
Method Object Indoor Outdoor TLS Map All Runtime (s)
edge graph edge graph edge graph edge graph edge graph edge graph
FPFH+FGR [zhou2016eccv] +PGO 9.7 2.1 1.2 0.4 2.2 0.5 2.7 0.6 4.3 1.6 2.5 0.7 19.7
FPFH+TEASER [yang2020tro] 14.8 10.2 5.1 2.3 9.5 3.0 13.8 7.9 26.1 9.8 10.6 5.9 22.2
KISS-Matcher [lim2025icra-kissmatcher] 16.4 13.6 15.9 5.7 21.8 7.4 51.6 25.1 75.1 63.5 34.4 17.3 18.4
Predator [huang2021cvpr-predator] 1.7 0.8 4.4 2.1 3.8 1.4 2.5 1.1 5.0 2.2 3.4 1.5 51.8
BUFFER [ao2023cvpr-buffer] 1.3 0.5 15.3 6.7 8.2 3.7 16.3 6.3 1.4 0.3 12.0 5.4 250.3
BUFFER-X [seo2025iccv] 46.1 36.5 14.2 3.4 28.3 15.0 42.2 20.4 11.5 2.9 28.9 13.7 36.0
BUFFER-X* [seo2025iccv] 52.8 40.1 21.4 7.7 36.9 21.2 49.1 25.8 17.3 6.6 36.1 20.8 36.0
SGHR [wang2023cvpr-sghr] 10.3 5.8 17.3 8.6 14.5 7.6 19.8 12.0 12.8 5.8 16.2 9.1 49.7
RPF [sun2025neurips] 1.8 1.1 0.2 0.1 0.3 0.1 0.4 0.1 1.0 0.6 0.5 0.2 11.5
RAP (Ours) 74.7 58.1 60.2 36.5 84.0 76.6 89.6 75.9 92.7 90.4 79.8 65.8 8.9
RAP (Ours) w/o RF 68.8 51.4 53.0 29.9 81.3 70.5 84.6 68.3 89.6 86.8 74.9 59.0 8.9
We additionally provide the ECDF plots for the rotation error, normalized translation error, and normalized global RMSE on the cross-domain multi-view registration benchmark in Fig.˜7 (overall) and Fig.˜8 (by scenario category). We compare RAP with the two best-performing baselines: BUFFER-X+PGO and KISS-Matcher+PGO. The results show that RAP outperforms the baselines across all metrics and scenarios consistently.
Figure 7: Comparison of the ECDF of the rotation error, normalized translation error, and normalized global RMSE on the cross-domain multi-view registration benchmark.
Figure 8: Comparison of the ECDF of the rotation error, normalized translation error, and normalized global RMSE on the five scenario categories of the cross-domain multi-view registration benchmark.
D.4 Pose estimation results for offline SLAM
One potential application of our model is offline SLAM, where we use it to estimate poses of LiDAR scans in a batch. We evaluate pose estimation accuracy using the absolute trajectory error (ATE) metric on the FusionPortablev2 dataset [wei2025ijrr-fusionportablev2], which is an unseen dataset for our model. We compare our model with state-of-the-art SLAM systems: R3LIVE [jiarong2022icra] (LiDAR + camera + IMU), FAST-LIO2 [wei2022tro] (LiDAR + IMU), VINS-Fusion [qin2018tro] (camera + IMU), and DROID-SLAM [teed2021neurips] (camera only) in Tab.˜9. We additionally show estimated trajectories of two sequences in Fig.˜10 and merged point cloud maps on FusionPortablev2 and the Newer College dataset (NCD) [ramezani2020iros] in Fig.˜9.
Although our model is trained with at most 16 views, it can handle an arbitrary number of views during inference since we do not condition on view indices. However, feeding thousands of LiDAR scans at once is computationally expensive and may cause out-of-memory issues. We therefore sample LiDAR scans at 2 s (20 frames) and 10 s (100 frames) for evaluation, resulting in hundreds or tens of frames per sequence. Note that our model does not assume sequential order of the LiDAR scans and does not rely on any initial pose guess, which are typically required by conventional SLAM systems.
As shown in Tab.˜9, our model achieves comparable or better accuracy than the SLAM baselines (R3LIVE, FAST-LIO2, VINS-Fusion) that fuse LiDAR, camera, and IMU data, especially on scenes with geometric degenerations (e.g., ugv_parking00). Intuitively, our model accomplishes an implicit LiDAR bundle adjustment, leading to better performance than approaches relying on incremental scan-to-map registration. The generation process takes about 30 s for more than 100 sampled frames (and more than 2000 frames in the original sequence) on an NVIDIA A5000 GPU, which is comparable to or faster than the compared baselines that process at about 30 ms per frame.
Table 9: SLAM localization accuracy comparison with the state-of-the-art SLAM systems on the FusionPortablev2 dataset [wei2025ijrr-fusionportablev2]. We calculate mean translation ATE [m] for each sequence. The best result is shown in bold, and the second best result is underlined. Our model works zero-shot on this dataset and takes the LiDAR point clouds per 2 s (20 frames) and 10 s (100 frames) in batches for evaluation. VINS-Fusion w/ LC means VINS-Fusion with loop closure enabled.
Method LiDAR Camera IMU handheld room00 handheld escalator00 legged grass00 legged room00 ugv parking00
R3LIVE [jiarong2022icra] 0.057 0.093 0.069 0.068 0.424
FAST-LIO2 [wei2022tro] $\times$ 0.058 0.085 0.327 0.093 0.271
VINS-Fusion w/ LC [qin2018tro] $\times$ 0.063 0.258 1.801 0.149 2.400
DROID-SLAM [teed2021neurips] $\times$ $\times$ 0.118 4.427 7.011 0.135 2.019
RAP (Ours) per 2 s $\times$ $\times$ 0.050 0.164 0.105 0.082 0.250
RAP (Ours) per 10 s $\times$ $\times$ 0.052 0.140 0.108 0.064 0.252
Figure 9: Offline SLAM results of our model on the FusionPortablev2 [wei2025ijrr-fusionportablev2] and the Newer College dataset [ramezani2020iros]. Our model works zero-shot on these datasets, taking LiDAR point clouds per 2 s (20 frames) in batches for evaluation. Different colors in the merged point cloud represent different LiDAR scans. We report the view count, generation runtime, and the localization ATE [m] for each sequence.
(a) handheld_escalator00 (XZ plane)
(b) ugv_parking00 (XY plane)
Figure 10: Estimated trajectories of RAP compared with the ground truth trajectories on the FusionPortablev2 dataset [wei2025ijrr-fusionportablev2]. Left: handheld_escalator00 sequence shown in the XZ plane. Right: ugv_parking00 sequence shown in the XY plane.
Appendix E Additional Information of the Cross-Domain Multi-View Registration Benchmark
Tab.˜10 provides an overview of the datasets used to compose the cross-domain multi-view registration benchmark separated by scenario category. Most of these datasets are originally not designed for registration evaluation. The detailed statistics of the datasets can be found in the overview figure (Fig.˜3).
Table 10: Overview of the datasets used to compose the cross-domain multi-view registration benchmark separated by scenario category.
Dataset Location Sensor License
Object
C3VD [bobrow2023c3vd] - clinical colonoscope CC BY-NC-SA 3.0
dSTORM nuclear lamina [vandelinde2011nprot-dstorm] - dSTORM microscope -
Fruit shape completion [magistri2025icra] Germany hand-held laser scanner, realsense RGB-D camera -
Stanford 3D Models [curless1996siggraph] USA Cyberware scanner Custom research license
Indoor Scan
3DCSR [huang2021arxiv-comprehensive] USA Kinect RGB-D camera, 16-beam LiDAR, SfM CC BY-NC-SA 3.0
Bonn RGB-D [palazzolo2019iros] Germany ASUS RGB-D camera -
IILABS3D [ribeiro2025access] Portugal Livox Mid-360, RoboSense Helios CC BY-SA
Leg-KILO [ou2024ral] China Livox Mid-360 CC BY 4.0
Matterport3D [chang2017threedv-matterport3d] USA Matterport RGB-D camera Custom research license
TIERS [sier2023rs-tiers] Finland Velodyne 16-beam, Ouster 64/128-beam, Livox LiDAR MIT
TUM RGB-D [sturm2012iros] Germany Kinect RGB-D camera CC BY 4.0
Outdoor Scan
Argoverse2 [wilson2021neurips] USA 2 x LiDARs CC BY-NC-SA 4.0
Digiforest [malladi2025icra] Switzerland Hesai 32/64-beam LiDAR -
MUST-C [chong2026sdata] Germany Ouster 128 beam LiDAR, RIEGL scanner, LMI scanner CC BY 4.0
NCLT [carlevaris-bianco2016ijrr] USA Velodyne 32-beam LiDAR ODbL
Nuscenes [caesar2020cvpr] USA, Singapore Velodyne 32-beam LiDAR CC BY-NC 4.0
PandaSet [xiao2021itsc-pandaset] USA Hesai 64-beam and foward-facing LiDAR CC BY-NC-SA 4.0
SGAB [chong2022sgab] Singapore 3 x Velodyne 16-beam LiDARs Custom research license
SubT [zhao2024cvpr-subt] USA Velodyne 16-beam LiDAR -
Truckscenes [fent2024neurips-truckscenes] Germany 6 x LiDARs CC BY-NC-SA 4.0
Waymo [sun2020cvpr] USA 5 x LiDARs Custom research license
ZOD [alibeigi2023iccv-zod] 14 European countries Velodyne 128-beam LiDAR CC BY-SA 4.0
TLS
ETH3D [schops2017cvpr-eth3d] Switzerland Leica TLS CC BY-NC-SA 4.0
ETH-TLS [theiler2015jprs] Switzerland TLS -
IndoorLRS [park2017iccv] - Faro TLS and Asus RGB-D camera -
MCD [nguyen2024cvpr-mcd] Sweden, Germany, Singapore Leica and Faro TLS CC BY-NC-SA 4.0
Oxford-TLS [tao2025ijrr-oxfordspires, ramezani2020iros] UK Leica TLS -
R3DS [robotic3dscanrepository] Germany, Croatia Riegl TLS -
RESSO [chen2020tgrs-plade] Netherlands, Saudi Arabia Leica and Faro TLS -
Semantic3D [hackel2017isprs] Switzerland TLS CC BY-NC-SA 3.0
Tongji-Trees [wang2023jprs-globalmatch] China Z+F TLS -
VMML [michailidis2019vc-aspire] Switzerland Faro TLS -
WHU-TLS [dong2020jprs-whutls] China RIEGL TLS -
Map
Abenberg [hebel2013jprs] Germany ALS CC BY-NC-SA 4.0
AHN [cserep2023jag-ahn] Netherlands ALS CC0
Alita [yin2022arxiv-alita] USA Map built by Velodyne 16-beam LiDAR BSD 3-Clause
Kimera-Multi [tian2023iros-kimeramultidata] USA Map built by Velodyne 16-beam LiDAR MIT
MS-HKUSTGZ [hu2024arxiv-msmapping] China Map built by HESAI 32-beam LiDAR, Leica TLS -
Appendix F Additional Qualitative Results
We provide additional qualitative results of our model on the pairwise and multi-view point cloud registration tasks on various datasets in this section, as shown in Fig.˜11, Fig.˜12, Fig.˜14, Fig.˜15, and Fig.˜13.
We also provide some failure cases of our model in Fig.˜16. It is shown that our model sometimes fails to correctly register the point clouds in the cases of very low overlap ratio or even no overlap due to the thickness of wall. From the examples shown for NSS dataset, it is shown that our model can still guess the reasonable global layout of the scene for the ambiguous failed cases.
Figure 11: Qualitative results of our model on pairwise registration with large scan interval (Waymo, KITTI, KAIST, NCD, MIT, TIERS) or low overlap ratio (NSS, 3DLoMatch).